OpenStudy (anonymous):

fFind the distance between the points given. (-3, 0) and (0,radical 7 ) do u handle it the same way? beacuase i keep on getting a diffrent anwser from the ones given

OpenStudy (campbell_st):

remember the index laws for power of a power \[(x^a)^b = x^{a \times b}\] this can be used here \[\sqrt{7} = 7^{\frac{1}{2}} ...then.... (7^{\frac{1}{2}})^2 = 7^{\frac{1}{2}\times2} = 7\] the distance is \[d = \sqrt{(-3 - 0)^2 + (0 - \sqrt{7})^2} = \sqrt{(-3)^2 + (\sqrt{7})^2} = \sqrt{9 + 7}\] so the distance is 4 units